OpenStudy (anonymous):

Solve the equation. Check the solution. -2/x+4 = 4/x+3

OpenStudy (anonymous):

1-Combine similar terms -2/x-4/x=3-4 Can you solve for x now ?

OpenStudy (anonymous):

can you show me?

OpenStudy (anonymous):

but that's not one of my choices?

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\frac{-2}{x+4}=\frac{4}{x+3}\iff -2(x+3)=4(x+4)\]

OpenStudy (anonymous):

i.e. cross multiply do you need help solving it now, or is it okay after this first step?

OpenStudy (anonymous):

it's okay I got it, thank you!!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Oh ,it wasn't clear ! next time draw it